OpenStudy (anonymous):

Solve: f(a+2)=2(a+2)^2-(a+2)+5

OpenStudy (anonymous):

I'd make (a+2) another variable then factor normally. Let (a+2) be k. f(k) = 2k^2 -k +5 -factor- k= something and something else something and something else = a+2, solve for a.